# Post-reduction inference for confidence sets of models

### v1

@@ 4.3. Co-sufficient information partitioning through randomisation
$f_{\bar{R}}(r)=\frac{\Gamma((n-d_{\theta})/2)}{\sqrt{\pi}\Gamma\{(n-d_{\theta}-1)/2\}}(1-r^{2})^{(n-d_{\theta}-3)/2},\quad-1<r<1.$ (4.9)
@@
[…]
@@
$(1-r^{2})^{(n-d_{\theta}-3)/2}\simeq e^{-(n-d_{\theta})r^{2}/2},\quad(r\rightarrow 0),$ (4.10)
@@
Conditional on |^$|\hat{\mathcal{E}}^{(1)}|=df_{1}$ and |^$|\hat{\mathcal{E}}^{(2)}|=df_{2}$, and on the containments ^$\mathcal{E}^{*}\subseteq\hat{\mathcal{E}}^{(1)}$ and ^$\mathcal{E}^{*}\subseteq\hat{\mathcal{E}}^{(2)}$,
@@
=$\stackrel{{\scriptstyle d}}{{=}}$ $\;R_{n}+O_{p}\Bigl{(}\max\Bigl{\{}k\nu^{-1/2}(n-d_{\theta})^{-1},k\nu^{-1}(n-d_{\theta})^{-1/2}\Bigr{\}}\Bigr{)},\quad(n,\nu\rightarrow\infty,),$
where $\nu=df_{1}+df_{2}\asymp n$. It follows from (4.7) that $\hat{R}_{n}\rightarrow_{d}N(0,1)$ provided that $k\lesssim n^{3/2}$.
@@
The error term is minimised with $k=2$ and in view of (4.10), we expect this choice to provide a close distributional approximation under the null model. There may be reasons to prefer larger $k$, particularly if this provides appreciable improvements in power to detect departures from the null model for only a small loss in coverage. Conditioning on |^$|\hat{\mathcal{E}}^{(1)}|$ and |^$|\hat{\mathcal{E}}^{(2)}|$ in the statement of Theorem 4.1 breaks the dependence between $\hat{Q}^{(1)},\ldots,\hat{Q}^{(k)}$. In the simulations of §5, where this aspect is implicitly averaged over in the simulation replicates, some mild dependence is introduced, which may contribute to the deterioration in the distribution theory for increasing $k$.
It may be questioned, in view of the discussion preceding Theorem 4.1, why the randomisation of equation (4.3) is needed at all, as rotational invariance of the Haar measure implies that any vector $v$ on a hypersphere in $\mathbb{R}^{n-d_{\theta}}$ gives rise to an angle $\langle Q,v\rangle$ that follows the distribution in (4.9). This approach, like that of §4.2 ignores the orientation of the embedding manifold $\mathcal{X}^{\perp}$ in $\mathbb{R}^{n}$. Thus, while the null distribution theory is correct in an unconditional sense, the analysis is not relevant for detecting departures from the null model.
@@ 4.4. A test based on the ancillary component
The result is a direct application of Polya’s theorem. An alternative derivation with possibly very conservative rates is given in Appendix B via a more explicit statement about the rate of convergence of $\hat{\sigma}^{2}$. ∎
4.5. On the combination of ancillary and co-sufficient information
In principle, a combination of the ancillary and co-sufficient components achieves the maximal extraction of information relevant for the assessment of model adequacy. In the Gaussian case under consideration here, and under the model postulated for a potential contradiction, a weak form of ancillarity for $\theta$ is used, in which dependence on the nuisance parameter $\sigma^{2}$ is allowed. This dependence, while less critical in conventional uses of ancillarity, is consequential for present purposes, as explained by the following observations.
The random vector $W$ contains both the co-sufficient and the ancillary information. Indeed, when $\sigma^{2}$ is known, there is no ancillary component and $W$ is co-sufficient. Thus, when $\sigma^{2}$ is known, tests of model adequacy would be based on $W$ via $\|W\|^{2}=R^{2}\stackrel{{\scriptstyle d}}{{=}}\sigma^{2}\chi^{2}_{n-d_{\theta}}$. The situation when $\sigma^{2}$ is unknown is more subtle. Although $R^{2}$ and $Q$ are independent, $R^{2}$ is not independent of $\tilde{Q}^{(1)},\ldots,\tilde{Q}^{(k)}$, the pseudo-replicates of $Q$ constructed with randomisation, which invalidates the use of Fisher’s combination test. Combination tests allowing for dependence are conservative, compromising the calibration accuracy that has been sought throughout the paper, and partially negating the effects of information recovery through combination. The situation is exacerbated once additional dependence is introduced through the estimation of $\sigma^{2}$.
@@ 5.1. Experimental design
This section explores the numerical performance of the approaches based on the co-sufficient and ancillary separations relative to the simpler approach based on sample splitting, which inevitably discards relevant information. We focus on the Gaussian linear model of §3.1. Conditional on all signal variables surviving the reduction phase, all three approaches have the correct theoretical coverage probability for sufficiently large sample size, but the approach based on the ancillary separation is expected to be more sensitive to estimation of the error variance in small samples, as discussed in §4.5. The extent of this effect is uncovered through simulation in idealised settings. The approach based on sample splitting discards information at both the reduction phase and the model assessment phase, and will therefore discard signal variables with higher probability in the initial phase for an encompassing set $\hat{\mathcal{E}}$ of given size, or conversely, include more variables in $\hat{\mathcal{E}}$ for a fixed probability $\mathop{\mathrm{pr}}(\mathcal{E}^{*}\subseteq\hat{\mathcal{E}})$.
@@
In 500 simulation replicates, $n$ independent rows of $X$ were generated from a normal distribution of dimension $p$ with zero mean and covariance matrix $\Sigma$ having a Toeplitz structure $\Sigma_{ij}=\rho^{|i-j|}$, for $\rho\in(0,1)$. The outcome $Y$ was generated as $Y\sim N(X\theta,\sigma^{2}I_{n})$, where $\sigma^{2}=1$ and $\theta=(t,t,t,0,\ldots,0)^{{\rm T}}$; in other words, the true model is =$\mathcal{E}^{*}=\{1,2,3\}$.
@@
Since the emphasis of this paper is on the model assessment phase, we simplify the reduction phase by choosing $p=400$, so that a single-stage version of Cox reduction can be used rather than the more general two-stage reductions that would be required with a considerably larger number of variables. If $p$ is larger, the reduction phase is more difficult, and the true set of signal variables are less likely in a reduced data set to survive reduction, strengthening the argument for a version that uses all data for both phases of analysis. Detailed analyses of Cox reduction, with extensive numerical comparisons for higher dimensions and different modelling frameworks, are available in Battey and Cox (2018), Hoeltgebaum and Battey (2019) and Lewis and Battey (2025). The latter paper provides insights into how the procedure used at the reduction phase interacts with the model assessment phase. Cox reduction for $p=400$ proceeds by arranging the variable indices at random in a $20\times 20$ grid and traversing the rows and columns, fitting a linear regression to each set of variables. Thus each variable is a covariate in two regressions, alongside different companion variables. Variables that are not significant at a given significance level in either analysis are discarded, the resulting set of variables forming the comprehensive model $\hat{\mathcal{E}}$. In a real application, the significance level in each 20-dimensional regression would ideally be chosen to give an acceptable level of stability for $\hat{\mathcal{E}}$ upon re-randomisation of the variable indices in the grid. For simplicity here, we start with a significance level of $0.05$, reducing it in increments of $0.001$ until the encompassing model $\hat{\mathcal{E}}$ contains at most 15 variables.
@@
Prior to the model-assessment phase, we estimated the error variance via $\mathop{\hat{\sigma}^{2}_{\mathrm{mrcv}}}$ as defined in (4.6), as both the co-sufficiency-based and ancillary-based tests require a consistent estimator of $\sigma^{2}$. This estimator was computed using the first $[0.6n]$ observations, in order to reduce the dependence between the estimator and the test statistics in the idealised known-variance setting. Estimation with the full sample led to invalid tests under the null, whereas the 60% version was found to work well over many scenarios. The conservative variable selection procedure employed for screening in the definition of $\mathop{\hat{\sigma}^{2}_{\mathrm{mrcv}}}$ was the undertuned lasso procedure described before. The lasso, having excellent predictive properties, delivers good estimates of variance even when signal variables are discarded in favour of other variables.
@@
  • Co-sufficient information: $k$ replicates $\hat{Q}_{1},\ldots,\hat{Q}_{k}$ of the spherical vector $Q$ were computed via randomisation as described in §4.3, and were tested for uniformity on the hypersphere in $n-d_{\theta}$ dimensions using the high-dimensional Rayleigh test.
@@
  • No adjustment for selection: for comparison, we compare the above constructions to the naïve approach of using the same data for reduction and for model assessment based on a likelihood ratio test of each submodel against the comprehensive model. Since we are in a Gaussian setting, the likelihood-ratio statistic reduces to the $F$ statistic. Under each null model $\mathcal{E}_{m}$ to be assessed, this $F$ statistic is calibrated against an $F$ distribution on |^$|\hat{\mathcal{E}}|-|\mathcal{E}_{m}|$ and $n-|\hat{\mathcal{E}}|$ degrees of freedom, which would be the exact distribution under the null model, had $\hat{\mathcal{E}}$ not been selected in the light of the data.
@@ 5.2. Coverage and expected size of the confidence sets
The quantities of interest are reported in Tables 44 for values of $n$, $t$ and $\rho$ each taken at one of two levels in combinations corresponding to a half-replicate of a factorial design: (low, low, high), (high, low, low), (low, high, low), (high, high, high). We also report in Table 5 part of the complementary half-replicate of the full factorial experiment, with each factor at its low level, as this illustrates more striking differences in coverage probabilities from using the co-sufficient or ancillary tests relative to the split-sample $F$ test. These differences are due to the low probability of survival of the full set of signal variables through the reduction phase.
The reported outcome variables are: the simulated probability that $\mathcal{E}^{*}$ is retained through the initial reduction; the simulated coverage probability of the confidence set $\mathcal{M}$; and the simulation average of the number of models in the confidence set. Numbers in parentheses indicate empirical standard deviations of the simulation averages.
@@
Cox + co-sufficient ($k=2$) 0.96 (0.01) 1.00 (0.00) 2402 (67)
Cox + co-sufficient ($k=8$) 0.88 (0.01) 1.00 (0.00) 1674 (60)
Cox + ancillary 0.86 (0.02) 1.00 (0.00) 1574 (59)
Cox + $F$ test 0.08 (0.01) 1.00 (0.00) 17 (2)
Lasso + co-sufficient ($k=2$) 0.94 (0.01) 0.99 (0.01) 2614 (61)
Lasso + co-sufficient ($k=8$) 0.86 (0.01) 0.99 (0.00) 1904 (57)
Lasso + ancillary 0.85 (0.02) 0.99 (0.00) 1821 (57)
@@
split Cox + split $F$ test 0.92 (0.01) 0.96 (0.01) 1209 (43)
split Lasso + split $F$ test 0.86 (0.02) 0.91 (0.01) 1441 (46)
Table 1. $n=100$; $t=0.5$; $\rho=0.5$.
@@
Cox + co-sufficient ($k=2$) 0.99 (0.01) 1.00 (0.00) 2624 (70)
Cox + co-sufficient ($k=8$) 0.96 (0.01) 1.00 (0.00) 1821 (69)
Cox + ancillary 0.95 (0.02) 1.00 (0.00) 1700 (68)
Cox + $F$ test 0.00 (0.00) 1.00 (0.00) 1 (0)
Lasso + co-sufficient ($k=2$) 0.99 (0.01) 1.00 (0.00) 2640 (67)
Lasso + co-sufficient ($k=8$) 0.95 (0.01) 1.00 (0.00) 1865 (67)
Lasso + ancillary 0.95 (0.02) 1.00 (0.00) 1749 (65)
@@
split Cox + split $F$ test 0.83 (0.02) 0.88 (0.01) 499 (27)
split Lasso + split $F$ test 0.91 (0.01) 0.97 (0.01) 490 (23)
Table 2. $n=150$; $t=0.5$; $\rho=0.1$.
@@
Cox + co-sufficient ($k=2$) 0.95 (0.01) 1.00 (0.00) 745 (51)
Cox + co-sufficient ($k=8$) 0.88 (0.01) 1.00 (0.00) 368 (34)
Cox + ancillary 0.87 (0.01) 1.00 (0.00) 334 (31)
Cox + $F$ test 0.36 (0.02) 1.00 (0.00) 27 (1)
Lasso + co-sufficient ($k=2$) 0.95 (0.01) 1.00 (0.00) 779 (47)
Lasso + co-sufficient ($k=8$) 0.88 (0.01) 1.00 (0.00) 393 (31)
Lasso + ancillary 0.88 (0.01) 1.00 (0.00) 362 (30)
@@
split Cox + split $F$ test 0.87 (0.01) 0.91 (0.01) 136 (10)
split Lasso + split $F$ test 0.96 (0.01) 0.97 (0.01) 132 (7)
Table 3. $n=100$; $t=1$; $\rho=0.1$.
@@
Cox + co-sufficient ($k=2$) 0.97 (0.01) 1.00 (0.00) 270 (13)
Cox + co-sufficient ($k=8$) 0.87 (0.01) 1.00 (0.00) 89 (5)
Cox + ancillary 0.86 (0.02) 1.00 (0.00) 80 (4)
Cox + $F$ test 0.62 (0.02) 1.00 (0.00) 44 (1)
Lasso + co-sufficient ($k=2$) 0.97 (0.01) 1.00 (0.00) 375 (14)
Lasso + co-sufficient ($k=8$) 0.89 (0.01) 1.00 (0.00) 113 (6)
Lasso + ancillary 0.86 (0.02) 1.00 (0.00) 98 (4)
@@
split Cox + split $F$ test 0.95 (0.01) 1.00 (0.00) 78 (3)
split Lasso + split $F$ test 0.95 (0.01) 1.00 (0.00) 86 (3)
Table 4. $n=150$; $t=1$; $\rho=0.5$.
Cox + co-sufficient ($k=2$) 0.91 (0.01) 0.92 (0.01) 3053 (70)
Cox + co-sufficient ($k=8$) 0.87 (0.02) 0.92 (0.01) 2477 (75)
Cox + ancillary 0.86 (0.02) 0.92 (0.01) 2396 (75)
Cox + $F$ test 0.02 (0.01) 0.92 (0.01) 4 (1)
Lasso + co-sufficient ($k=2$) 0.97 (0.01) 0.98 (0.01) 3347 (66)
Lasso + co-sufficient ($k=8$) 0.93 (0.01) 0.98 (0.01) 2750 (75)
Lasso + ancillary 0.92 (0.01) 0.98 (0.01) 2681 (76)
Lasso + $F$ test 0.00 (0.00) 0.98 (0.01) 0 (0)
split Cox + split $F$ test 0.49 (0.02) 0.51 (0.02) 1379 (59)
split Lasso + split $F$ test 0.75 (0.02) 0.79 (0.02) 1532 (62)
Table 5. $n=100$; $t=0.5$; $\rho=0.1$.
The half-replicate allows estimation of the marginal effects of increasing each of $n$, $t$ and $\rho$ from its low to its high level. These are presented in Table 6, where the reported effect on the coverage probability is multiplicative on the odds scale, with values close to one indicative of a null effect, and the reported effect on the expected size is additive on the log scale, with values close to zero suggesting a null effect. Some caution is needed in interpretation of the multiplicative marginal effects, as logistic regression is non-collapsible.
@@
Cox + co-sufficient ($k=2$) 2.65 0.51 0.64 -0.46 -1.72 -0.55
Cox + co-sufficient ($k=8$) 1.73 0.53 0.53 -0.67 -2.27 -0.75
Cox + ancillary 1.68 0.57 0.55 -0.68 -2.30 -0.75
Cox + $F$ test -1.17 2.12 1.66
Lasso + co-sufficient ($k=2$) 3.28 0.63 0.52 -0.36 -1.58 -0.37
Lasso + co-sufficient ($k=8$) 1.85 0.71 0.60 -0.63 -2.19 -0.61
Lasso + ancillary 1.68 0.65 0.50 -0.67 -2.25 -0.63
@@
split Cox + split $F$ test 1.10 1.53 2.59 -0.72 -2.02 0.16
split Lasso + split $F$ test 1.76 2.71 0.69 -0.75 -2.06 0.33
Table 6. Point estimates of marginal effects of $n$, $t$ and $\rho$.

### v2

@@ 5.2. Co-sufficient information partitioning through randomisation
(5.7) $f_{\bar{R}}(r)=\frac{\Gamma((n-d_{\theta})/2)}{\sqrt{\pi}\Gamma\{(n-d_{\theta}-1)/2\}}(1-r^{2})^{(n-d_{\theta}-3)/2},\quad-1<r<1.$
@@
$\text{var}(\bar{R})=$ $\;\frac{\Gamma(\tfrac{n-d_{\theta}}{2})}{\sqrt{\pi}\Gamma(\tfrac{n-d_{\theta}-1}{2})}\int_{0}^{1}s^{1/2}(1-s)^{(n-d_{\theta}-3)/2}ds$
@@
(5.8) $(1-r^{2})^{(n-d_{\theta}-3)/2}\simeq e^{-(n-d_{\theta})r^{2}/2},\quad(r\rightarrow 0),$
@@
Under the conditions of Proposition 5.2,
@@
$\stackrel{{\scriptstyle d}}{{=}}$ $\;R_{n}+O_{p}\bigl(\max\{k\nu^{-3/2},n^{-1/2}\}\bigr),\quad(n,\nu\rightarrow\infty,),$
where $\nu=df_{1}+df_{2}\asymp n$. It follows from (5.5) that $\hat{R}_{n}\rightarrow_{d}N(0,1)$ provided that $k\lesssim n^{1/2}$.
@@
The error term is minimised with $k=2$ and, in view of (5.8), we expect this choice to provide a close distributional approximation under the null model. There may be reasons to prefer larger $k$, particularly if this provides appreciable improvements in power to detect departures from the null model for only a small loss in coverage. Conditioning on $|\hat{\mathcal{E}}^{(1)}|$ and $|\hat{\mathcal{E}}^{(2)}|$ in the statement of Theorem 5.1 breaks the dependence between $\hat{Q}^{(1)},\ldots,\hat{Q}^{(k)}$. In the simulations of §6, where this aspect is implicitly averaged over in the simulation replicates, some mild dependence is introduced, which may contribute to the deterioration in the distribution theory for increasing $k$.
It may be questioned, in view of the discussion preceding Theorem 5.1, why the randomisation of equation (5.1) is needed at all, as rotational invariance of the Haar measure implies that any vector $v$ on a hypersphere in $\mathbb{R}^{n-d_{\theta}}$ gives rise to an angle $\langle Q,v\rangle$ that follows the distribution in (5.7). Such an approach, like that of § A ignores the orientation of the embedding manifold $\mathcal{X}^{\perp}$ in $\mathbb{R}^{n}$. Thus, while the null distribution theory is correct in an unconditional sense, the analysis is not relevant for detecting departures from the null model.
@@ 5.3. A test based on the ancillary component
The result is a direct application of Polya’s theorem. An alternative derivation with possibly very conservative rates is given in Appendix C via a more explicit statement about the rate of convergence of $\hat{\sigma}^{2}$. ∎
@@ 6.1. Experimental design
This section explores the numerical performance of the approaches based on the co-sufficient and ancillary separations relative to the simpler approach based on sample splitting, which inevitably discards relevant information. We focus on the Gaussian linear model of §4.1. Conditional on all signal variables surviving the reduction phase, all three approaches have the correct theoretical coverage probability for sufficiently large sample size, but the approach based on the ancillary separation is expected to be more sensitive to estimation of the error variance in small samples, as discussed at the end of the previous section. The extent of this effect is uncovered through simulation in idealised settings. The approach based on sample splitting discards information at both the reduction phase and the model assessment phase, and in small samples will therefore discard signal variables with higher probability in the initial phase for an encompassing set $\hat{\mathcal{E}}$ of given size, or conversely, include more variables in $\hat{\mathcal{E}}$ for a fixed probability $\mathop{\mathrm{pr}}(\mathcal{E}^{*}\subseteq\hat{\mathcal{E}})$.
@@
In 500 simulation replicates, $n$ independent rows of $X$ were generated from a normal distribution of dimension $p$ with zero mean and covariance matrix $\Sigma$, where $\Sigma$ is the identity matrix with the top-left $10\times 10$ block replaced by an equi-correlation matrix with correlation $\rho\in(0,1)$. The outcome $Y$ was generated as $Y\sim N(X\theta,\sigma^{2}I_{n})$, where $\sigma^{2}=1$ and $\theta=(t,t,t,0,\ldots,0)^{{\rm T}}$; in other words, the true model is $\mathcal{E}^{*}=\{1,2,3\}$. The three signal variables were therefore correlated with each other and with 7 irrelevant variables.
@@
Since the emphasis of this paper is on the model assessment phase, we simplify the reduction phase by choosing $p=400$, so that a single-stage version of Cox reduction can be used rather than the more general two-stage reductions that would be required with a considerably larger number of variables. If $p$ is larger, the reduction phase is more difficult, and the true set of signal variables are less likely in a reduced data set to survive reduction, strengthening the argument for a version that uses all data for both phases of analysis. Detailed analyses of Cox reduction, with extensive numerical comparisons for higher dimensions and different modelling frameworks, are available in Battey and Cox (2018), Hoeltgebaum and Battey (2019) and Lewis and Battey (2025). The latter paper provides insights into how the procedure used at the reduction phase interacts with the model assessment phase. Cox reduction for $p=400$ proceeds by arranging the variable indices at random in a $20\times 20$ grid and traversing the rows and columns, fitting a linear regression to each set of variables. Thus each variable is a covariate in two regressions, alongside different companion variables. Variables that are not significant at a given significance level in either analysis are discarded, the resulting set of variables forming the comprehensive model $\hat{\mathcal{E}}$. In a real application, the significance level in each 20-dimensional regression would ideally be chosen to give an acceptable level of stability for $\hat{\mathcal{E}}$ upon re-randomisation of the variable indices in the grid. For simplicity here, we start with a significance level of $0.05$, reducing it in increments of $0.001$ until the encompassing model $\hat{\mathcal{E}}$ contains at most 15 variables. In very high-dimensional problems, a provisional reduction to 15 variables is likely to be too severe; the number can in principle be any value smaller than $n$, the main difficulty being the computational burden associated with testing a large number of submodels. This, however, is less problematic in a single application than in the many-replicate simulation considered here.
@@
Prior to the model-assessment phase, we estimated the error variance via $\mathop{\hat{\sigma}^{2}_{\mathrm{mrcv}}}$ as defined in (5.4), as both the co-sufficiency-based and ancillary-based tests require a consistent estimator of $\sigma^{2}$. This estimator was computed using the first $[0.6n]$ observations, in order to reduce the dependence between the estimator and the test statistics in the idealised known-variance setting. Estimation with the full sample led to invalid tests under the null, whereas the 60% version worked well over many scenarios. The conservative variable selection procedure employed for screening in the definition of $\mathop{\hat{\sigma}^{2}_{\mathrm{mrcv}}}$ was the undertuned lasso procedure described before. The lasso, having excellent predictive properties, delivers good estimates of variance even when signal variables are discarded in favour of other variables.
@@
  • Co-sufficient information: $k$ replicates $\hat{Q}_{1},\ldots,\hat{Q}_{k}$ of $Q$ were computed as described in §5.2, and were tested for uniformity on the hypersphere in $n-d_{\theta}$ dimensions using the high-dimensional Rayleigh test.
@@
  • No adjustment for selection: for comparison, we compare the above constructions to the naïve approach of using the same data for reduction and for model assessment based on a likelihood ratio test of each submodel against the comprehensive model. Since we are in a Gaussian setting, the likelihood-ratio statistic reduces to the $F$ statistic. Under each null model $\mathcal{E}_{m}$ to be assessed, this $F$ statistic is calibrated against an $F$ distribution on $|\hat{\mathcal{E}}|-|\mathcal{E}_{m}|$ and n|^|n-$\hat{\mathcal{E}}|$ degrees of freedom, which would be the exact distribution under the null model, had $\hat{\mathcal{E}}$ not been selected in the light of the data.
@@ 6.2. Coverage and expected size of the confidence sets
The quantities of interest are reported in Tables 1–4 for values of $n$, $t$ and $\rho$ each taken at one of two levels in combinations corresponding to a half-replicate of a factorial design: (low, low, low), (high, high, low), (high, low, high), (low, high, high). The reported outcome variables are: the simulated probability that $\mathcal{E}^{*}$ is retained through the initial reduction; the simulated coverage probability of the confidence set $\mathcal{M}$ at nominal coverage 0.95; and the simulation average of the number of models in the confidence set. Numbers in parentheses indicate empirical standard deviations of the simulation averages.
Table 1. $n=100$; $t=0.5$; $\rho=0.1$.
@@
Cox + co-sufficient ($k=2$) 0.94 (0.02) 0.94 (0.02) 3097 (147)
Cox + co-sufficient ($k=8$) 0.91 (0.03) 0.94 (0.02) 2399 (169)
Cox + ancillary 0.90 (0.03) 0.94 (0.02) 2320 (166)
Cox + $F$ test 0.02 (0.01) 0.94 (0.02) 4 (1)
Lasso + co-sufficient ($k=2$) 0.98 (0.01) 0.99 (0.01) 3455 (149)
Lasso + co-sufficient ($k=8$) 0.96 (0.02) 0.99 (0.01) 2763 (161)
Lasso + ancillary 0.95 (0.02) 0.99 (0.01) 2651 (160)
@@
split Cox + split $F$ test 0.49 (0.05) 0.51 (0.05) 1485 (129)
split Lasso + split $F$ test 0.76 (0.04) 0.78 (0.04) 1485 (133)
Table 2. $n=120$; $t=1$; $\rho=0.1$.
@@
Cox + co-sufficient ($k=2$) 0.96 (0.01) 1.00 (0.00) 332 (24)
Cox + co-sufficient ($k=8$) 0.88 (0.01) 1.00 (0.00) 137 (13)
Cox + ancillary 0.86 (0.02) 1.00 (0.00) 124 (12)
Cox + $F$ test 0.41 (0.02) 1.00 (0.00) 31 (1)
Lasso + co-sufficient ($k=2$) 0.96 (0.01) 1.00 (0.00) 402 (29)
Lasso + co-sufficient ($k=8$) 0.88 (0.01) 1.00 (0.00) 158 (14)
Lasso + ancillary 0.86 (0.02) 1.00 (0.00) 146 (13)
@@
split Cox + split $F$ test 0.94 (0.01) 0.99 (0.00) 77 (3)
split Lasso + split $F$ test 0.95 (0.01) 1.00 (0.00) 84 (3)
Table 3. $n=120$; $t=0.5$; $\rho=0.5$.
@@
Cox + co-sufficient ($k=2$) 0.97 (0.01) 1.00 (0.00) 2863 (67)
Cox + co-sufficient ($k=8$) 0.92 (0.01) 1.00 (0.00) 1872 (65)
Cox + ancillary 0.90 (0.01) 1.00 (0.00) 1747 (64)
Cox + $F$ test 0.37 (0.02) 1.00 (0.00) 54 (3)
Lasso + co-sufficient ($k=2$) 0.98 (0.01) 1.00 (0.00) 2597 (55)
Lasso + co-sufficient ($k=8$) 0.91 (0.01) 1.00 (0.00) 1841 (54)
Lasso + ancillary 0.90 (0.01) 1.00 (0.00) 1745 (53)
@@
split Cox + split $F$ test 0.93 (0.01) 0.98 (0.01) 1200 (44)
split Lasso + split $F$ test 0.91 (0.01) 0.95 (0.01) 1121 (37)
Table 4. $n=100$; $t=1$; $\rho=0.5$.
@@
Cox + co-sufficient ($k=2$) 0.94 (0.01) 1.00 (0.00) 620 (29)
Cox + co-sufficient ($k=8$) 0.86 (0.02) 1.00 (0.00) 228 (13)
Cox + ancillary 0.83 (0.02) 1.00 (0.00) 200 (11)
Cox + $F$ test 0.74 (0.02) 1.00 (0.00) 53 (1)
Lasso + co-sufficient ($k=2$) 0.95 (0.01) 1.00 (0.00) 642 (27)
Lasso + co-sufficient ($k=8$) 0.85 (0.02) 1.00 (0.00) 305 (16)
Lasso + ancillary 0.83 (0.02) 1.00 (0.00) 265 (14)
@@
split Cox + split $F$ test 0.96 (0.01) 1.00 (0.00) 248 (11)
split Lasso + split $F$ test 0.96 (0.01) 1.00 (0.00) 237 (10)
The half-replicate of Tables 1–4 allows estimation of the main effects of increasing each of $n$, $t$ and $\rho$ from its low to its high level. These are presented in Table 5, where the reported effect on the coverage probability is multiplicative on the odds scale, with values close to one indicative of a null effect, and the reported effect on the expected size is additive on the log scale, with values close to zero suggesting a null effect.
Table 5. Point estimates of effects of $n$, $t$ and $\rho$ from the half-replicate of the factorial design (Tables 1–4). Reported effects are multiplicative on the odds scale for $\mathbb{P}(\mathcal{E}^{*}\subseteq\mathcal{M})$ and additive on the log scale for $\mathbb{E}(|\mathcal{M}|)$.
@@
Cox + co-sufficient ($k=2$) 1.334 0.929 1.077 -0.176 -1.250 0.005
Cox + co-sufficient ($k=8$) 1.080 0.789 0.988 -0.064 -1.133 -0.052
Cox + ancillary 1.059 0.780 0.945 -0.076 -1.116 -0.062
Cox + $F$ test 1.628 3.582 3.294 0.136 0.140 0.311
Lasso + co-sufficient ($k=2$) 1.061 0.660 0.943 -0.088 -1.264 -0.053
Lasso + co-sufficient ($k=8$) 0.859 0.643 0.755 -0.110 -1.184 -0.089
Lasso + ancillary 0.878 0.647 0.784 -0.117 -1.164 -0.097
@@
split Cox + split $F$ test 1.733 2.330 2.147 -0.157 -0.776 -0.040
split Lasso + split $F$ test 1.261 1.943 1.418 -0.171 -0.735 -0.071